OpenStudy (anonymous):

Consider the functions f(x) = 3x^2 and g(x) = −2x − 5. What is the value of f[g(−4)]?

OpenStudy (campbell_st):

1st find g(-4) g(-4) = -2(-4) - 5 g(-4) = 3 now find f(3) f(3) = 3(3)^2 = 27

OpenStudy (anonymous):

thanks